# C-POD: An AWS Cloud Framework for Edge Pod Automation and Remote Wireless Testbed Sharing

### v1

@@ I Introduction
OTA experiments on testbeds are a key step in bridging the gap between theory and practice in wireless research [7]. These experiments enable researchers to validate models, identify implementation challenges, and refine designs in controlled but realistic environments. In recent years, significant efforts have been made by the wireless research community toward the development of such platforms. Notable examples include the NSF PAWR platforms POWDER [breen2020powder], COSMOS [rauchaudhari2020cosmos], AERPAW [marojevic2020aerpaw], and ARA [zhang22ara], Colosseum [bonati2021colosseum], and the NextGen end-to-end network testbed [3], among others.
@@
To bridge this gap, we introduce C-POD, a cloud-native framework that automates the deployment and management of edge pods for remote wireless testbed access. Using a combination of cloud and edge resources, C-POD reduces the barrier to OTA experimentation by enabling researchers to share and access geographically distributed testbeds without extensive local infrastructure. It is particularly beneficial for small research groups that lack the resources or expertise to develop and maintain dedicated web portals and testbed management infrastructure. With its ready-to-use cloud native toolkit, C-POD makes these testbeds remotely accessible and easily shareable, enhancing their utility and impact. Moreover, C-POD supports dynamic edge scaling, persistent remote sessions, and stable graphical interfaces, making it well suited for long-running and high-variability experiments.
@@
  • Cloud-Enabled Edge Pod Automation Framework: We design a cloud-native, end-to-end framework based on AWS [aws] and Rancher Kubernetes Engine v2 (RKE2) [8] to provide isolated, resource-capped experiment pods on demand at the testbed edge.
@@ II C-POD Framework and Toolkit Design
Upon receiving user requests, a key function of the Federation Plane is to initiate the appropriate service automation workflows. This process is managed by the Function Manager Module, which interacts with all other components of the plane, such as the storage repository, database engine, and operations manager, through dedicated APIs. For example, when a user request involves accessing the Storage Repository, the Function Manager coordinates file transfers by utilizing temporary storage for fast data movement and community libraries for commonly used code. If the services initiated by the Function Manager require container images, the Image Repository is invoked to fetch the requested container script images.
Finally, the Operations Manager Module orchestrates interactions with edge devices within the Testbed Plane by deploying the SSM agent on the testbed’s control plane and maintaining continuous communication with it. This module also coordinates with the elastic VM to launch the noVNC client for browser-based remote access, leveraging the Cache Operator to configure port mappings and establish reverse SSH tunnels to the testbed. This deployment process ensures a seamless transition from the cloud-native Federation Plane to the edge resources.
@@ II-A Toolkit Development
From that point on, Kubelet continuously performs readiness and liveness probes, reports pod and node status back to the control plane, and takes corrective actions such as restarting failed pods or managing resources under low-capacity conditions. Once a container is created, the Kubelet configures the probes and resource metrics reporting, after which the container runtime executes a custom startup script. This script establishes a reverse SSH tunnel back to an Elastic Virtual Machine (VM), starts TurboVNC on the node, and launches XFCE4 under a lightweight Xvfb server. On the VM side, a web service based on noVNC listens on the corresponding port, connects through the SSH tunnel to localhost, and renders the XFCE4 desktop of the container in the browser of the experimenter.
@@ III-B2 Remote Access to SDR Testbed and Experiment
During the reserved time slot, the user navigates to the experiment section in the sidebar, where a Connect button becomes available. Clicking Connect opens a modal dialog listing all active session instances and offering the option Open New Instance. Selecting this option will launch the remote experiment interface, which presents a virtual desktop environment equipped with an embedded terminal for running experiments. Then, the experimenter can initiate their experiments in the remote terminal.
@@ III-D Cloud-Edge Latency Measurement
Specifically, we generate a continuous 10 Mbps UDP stream from the AWS cloud to the edge pod using the open source network performance tool iperf3. Packet arrival timestamps are captured at the edge with the command line packet analyzer tcpdump to display the network traffic, and the traces are analyzed offline to compute inter-arrival jitter. As shown in Fig. 8, the observed jitter remains consistently low, typically ranging from 0.3 ms to 0.8 ms, with only two spikes exceeding 1.75 ms around the 4-5 s and 36-37 s marks. The overall average jitter was approximately 0.6 ms, with the 95th percentile staying below 1.0 ms. These results indicate that packet arrival times varied slightly, suggesting that C-POD is able to process incoming UDP flows with very little buffering. Even under sustained 10 Mbps UDP load, the network path maintains sub-millisecond variability for the vast majority of the test duration. This low and predictable jitter profile demonstrates that the edge-in-pod deployment provides reliable, low-latency connectivity to the Federation Plane, which is essential for OTA experiments supporting next generation wireless networks.
Figure 8: Latency between Federation Plane and deployed pods.
@@ IV Conclusions
In this work, we have presented C-POD, a cloud-powered framework that automates the deployment and management of edge pods to enable scalable and accessible OTA experimentation on heterogeneous wireless testbeds. Using public cloud resources and distributed edge infrastructure, C-POD lowers the barrier to OTA experimentation and enables seamless testbed sharing without requiring extensive local setup. A prototype deployed on AWS and integrated with RF sensing and 5G testbeds demonstrates consistent resource utilization, uniform pod startup behavior, and reliable horizontal scaling. In future work, we will enhance C-POD’s elasticity through dynamic scaling, minimize pod initialization latency, and expand automation to more effectively support heterogeneous resources across multiple testbeds.

### v2

@@ I Introduction
OTA experiments on testbeds are a key step in bridging the gap between theory and practice in wireless research. These experiments enable researchers to validate models, identify implementation challenges, and refine designs in controlled but realistic environments. In recent years, significant efforts have been made by the wireless research community toward the development of such platforms. Notable examples include the NSF PAWR platforms POWDER [breen2020powder], COSMOS [rauchaudhari2020cosmos], AERPAW [marojevic2020aerpaw], and ARA [zhang22ara], Colosseum [bonati2021colosseum], and the NextGen end-to-end network testbed [3], among others.
@@
To bridge this gap, built on our previous work [CloudRAFT, NEXTCOMCOM, UnionLabs, WiNTECH24], in this paper we further introduce C-POD, a cloud-native framework that automates the deployment and management of edge pods for remote wireless testbed access. Using a combination of clod and edge resources, C-POD reduces the barrier to OTA experimentation by enabling researchers to share and access geographically distributed testbeds without extensive local infrastructure. It is particularly beneficial for small research groups that lack the resources or expertise to develop and maintain dedicated web portals and testbed management infrastructure. With its ready-to-use cloud native toolkit, C-POD makes these testbeds remotely accessible and easily shareable, enhancing their utility and impact. Moreover, C-POD supports dynamic edge scaling, persistent remote sessions, and stable graphical interfaces, making it well suited for long-running and high-variability experiments.
@@
  • Cloud-Enabled Edge Pod Automation Framework: We design a cloud-native, end-to-end framework based on Amazon Web Service (AWS) and Rancher Kubernetes Engine v2 (RKE2) [7] to provide isolated, resource-capped experiment pods on demand at the testbed edge.
@@ II C-POD Framework and Toolkit Design
Upon receiving user requests, a key function of the Federation Plane is to initiate the appropriate service automation workflows. This process is managed by the Function Manager Module, which interacts with all other components of the plane, such as the storage repository, database engine, and operations manager, through dedicated APIs. For example, when a user request involves accessing the Storage Repository, the Function Manager coordinates file transfers by utilizing temporary storage for fast data movement and community libraries for commonly used code. If the services initiated by the Function Manager require container images, the Image Repository is invoked to fetch the requested container images.
Finally, the Operations Manager Module orchestrates interactions with edge devices within the Testbed Plane by deploying the SSM agent on the testbed’s control plane and maintaining continuous communication with it. This module also coordinates with the elastic VM to launch the noVNC client for browser-based remote access, leveraging the Cache Operator to configure port mappings and establish reverse SSH tunnels to the testbed. This ensures a seamless transition from the cloud-native Federation Plane to the edge resources.
@@ II-A Toolkit Development
Then, Kubelet continuously performs readiness and liveness probes, reports pod and node status back to the control plane, and takes corrective actions such as restarting failed pods or managing resources under low-capacity conditions. Once a container is created, the Kubelet configures the probes and resource metrics reporting, after which the container runtime executes a custom startup script. This script establishes a reverse SSH tunnel back to an Elastic Virtual Machine (VM), starts TurboVNC on the node, and launches XFCE4 under a lightweight Xvfb server. On the VM side, a web service based on noVNC listens on the corresponding port, connects through the SSH tunnel to localhost, and renders the XFCE4 desktop of the container in the experimenter browser.
@@ III-B2 Remote Access to SDR Testbed and Experiment
During the reserved time slot, the user navigates to the experiment section in the sidebar, where a Connect button becomes available. Clicking Connect opens a modal dialog listing all active session instances and offering the option Open New Instance. Selecting this option will launch the remote experiment interface, which presents a virtual desktop environment equipped with an embedded terminal for running experiments. Then, the experimenter can initiate their experiments there.
@@ III-B3 Link Establishment on Remote OAI 5G Testbed
Figure 8: Latency between Federation Plane and deployed pods.
@@ III-D Cloud-Edge Latency Measurement
Specifically, we generate a continuous 10 Mbps UDP stream from the AWS cloud to the edge pod using the open source network performance tool iperf3. Packet arrival timestamps are captured at the edge with the command line packet analyzer tcpdump to display the network traffic, and the traces are analyzed offline to compute inter-arrival jitter. As shown in Fig. 8, the observed jitter remains consistently low, typically ranging from 0.3 ms to 0.8 ms, with only two spikes exceeding 1.75 ms around the 4-5 s and 36-37 s marks. The overall average jitter was approximately 0.6 ms, with the 95th percentile staying below 1.0 ms. These results indicate that packet arrival times varied slightly, suggesting that C-POD is able to process incoming UDP flows with very little buffering. Even under sustained 10 Mbps UDP load, the network path maintains sub-millisecond variability for the vast majority of the test duration. This low and predictable jitter profile demonstrates that the edge-in-pod deployment provides reliable, low-latency connectivity to the Federation Plane.
@@ IV Conclusions
In this work, we have presented C-POD, a cloud-powered framework that automates the deployment and management of edge pods to enable scalable and accessible OTA experimentation on heterogeneous wireless testbeds. Using public cloud resources and distributed edge infrastructure, C-POD lowers the barrier to OTA experimentation and enables seamless testbed sharing without requiring extensive local setup. A prototype deployed on AWS and integrated with RF sensing and 5G testbeds demonstrates consistent resource utilization, uniform pod startup behavior, and reliable horizontal scaling.